ethan ran two and a half times as far as conor. will ran 3 more miles than conor. together they ran a total of 27 miles. how far dideach run?
Answer by faceoff57(108)

e = 5/2c
w = 3 + c
c
5/2c + (3+c) + c = 27
5/2c + 3 + c + c = 27
5/2c + 2c + 3 = 27
5/2c + 4/2c + 3 = 27
9/2c = 27 - 3
9/2c = 24
9c = 48
c = 48/9
conor = 16/3 or 5 1/3 miles
ethan = (5/2)(16/3)
ethan = 40/3 or 13 1/3 miles
will = 8 1/3 miles